Shift report 20030208 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Feb 08 2003 - 08:02:13 EST

  • Next message: Apache: "Shift report 20030208 08:00-16:00"
    ------------------------------------------------------------
    Date: Sat Feb  8 08:02:13 2003, Shift: 00:00-08:00
    Supervisor: kh, with: 
    Log book#14 pages 5 - 5
    
    Shift Summary:
    ------------------------------------------------------------
    Beam from 3:00 through 3:45
    and 7:00 through shift end
    Runs  trig 3   trig 6
    7836   57K      46K
    7837   51K      20K
    move MRS to 40deg and change D5 to 1050A
    7838   in progress
    
    Plan for next period of time:
    MRS:
    500K at current setting (40 deg D5=1050A)
    500K at 40 deg D5=1050B
    
    FS
    130K at current setting (8 deg 1/4B)
    move FS to 12 deg
    200K at 12 deg 1/8B
    200K at 12 deg 1/20B
    150K at 12 deg 1/4B
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Feb 08 2003 - 08:02:58 EST